the sum of the two numbers is 39. When the smaller number is doubled and the larger is increased by 3, the resulting numbers are equal. What are the numbers?
Answer by richwmiller(17219)

a+b=39
2a=b+3
a=39-b
substitute 39-b for a in equation #2